A student is pushing a 70kg box across a floor by applying a 363N force to the right. The coefficient of static friction between the box and the floor is 0.5 and the coefficient of kinetic friction is 0.43. Neglect drag.
How many forces are acting on the box?

What direction is the friction on the box?

Is the friction static or kinetic?

What is the magnitude of frictional force on the box?

What is the magnitude of the net force on the box?

What is the magnitude of box's acceleration?
1
Expert's answer
2020-05-05T18:42:04-0400

1) 4

2) Opposite motion.

3) Kinetic.

4)


"F_f=\mu m g=(0.43)(70)(9.8)=295\ N"

5)


"F_{net}=363-295=68\ N"

6)


"a=\frac{68}{70}=0.97\frac{m}{s^2}"